NBA players ~ Michael wants to investigate whether the distribution of his classmates’ favorite NBA players of all time is different from the one reported by NBA. He randomly surveys his classmates. The table below contains the percentages reported by NBA and the observed counts from his random sample.

Player

LeBron James

Kobe Bryant

Steph Curry

Michael Jordan

Others

NBA

Percentages

20%

30%

15%

19%

16%

Observed

Counts

26

28

10

15

11

Michael conducts the appropriate hypothesis test.

If, in fact, the percentages are as reported by NBA, then what is the contribution of LeBron James category to Michael's observed test statistic?
Answer #1

Total Observed Counts = 26+28+10+15+11 = 90

LeBron James's observed count = 26

LeBron James's expected value = 90×0.20 = 18

The contribution of LeBron James category to Michael's observed test statistic

= (Observed - Expected)^2/Expected

= (26 - 18)^2/18

= 3.555555

= 3.5556 (to four decimal places)